OpenStudy (anonymous):

A cylinder has a volume of 15 cubic inches. What is the volume of a cone that fits exactly inside the cylinder?

OpenStudy (mathstudent55):

Do you know the formulas for the volumes of a cylinder and a cone? Look them up. You'll find the answer to the question easily if you see the formulas.

OpenStudy (anonymous):

1/3(15inches^3) = ? not really sure what to do here

OpenStudy (anonymous):

thats what it says in my lesson anyways

OpenStudy (mathstudent55):

Forget the numbers. Just write the formulas the way they'd be written in a textbook.

OpenStudy (anonymous):

I don't have a textbook :/

OpenStudy (mathstudent55):

Oh, wait. You already answered correctly above.

OpenStudy (anonymous):

I'm stuck though. I'm not sure how to solve it?

OpenStudy (mathstudent55):

Ok, I'll be your textbook for this problem. Here are the formulas you need for this problem. \(\large V_{cylinder} = \pi r^2 h\) \(\large V_{cone} = \dfrac{1}{3}\pi r^2 h\)

OpenStudy (anonymous):

it says "math processing error"

OpenStudy (mathstudent55):

Notice both formulas a very similar. The both have the part \(\pi r^2 h\) What is the difference between them?

OpenStudy (anonymous):

what formulas did you type? I'll look them up quick cause it won't show me

OpenStudy (mathstudent55):